ten years from now, orlando will be three times older than he is today. what is his current age word problems
Answer by srinivas.g(540)

let x be the age of Orlando now .
after 10 years from now his age will be x+10
but it is mentioned that his age is 3 times older than present age
hence x+10 = 3x
subtract x on both sides
x+10-x = 3x-x
10 =2x
divide with 2 on both sides
10/2 = 2x/2
5 =x
hence current age of Orlando is 5 years